# RIGNO: A Graph-based framework for robust and accurate operator learning for PDEs on arbitrary domains

### v1

@@ Abstract
Learning the solution operators of PDEs on arbitrary domains is challenging due to the diversity of possible domain shapes, in addition to the often intricate underlying physics. We propose an end-to-end graph neural network (GNN) based neural operator to learn PDE solution operators from data on point clouds in arbitrary domains. Our multi-scale model maps data between input/output point clouds by passing it through a downsampled regional mesh. Many novel elements are also incorporated to ensure resolution invariance and temporal continuity. Our model, termed RIGNO, is tested on a challenging suite of benchmarks, composed of various time-dependent and steady PDEs defined on a diverse set of domains. We demonstrate that RIGNO is significantly more accurate than neural operator baselines and robustly generalizes to unseen spatial resolutions and time instances.
@@ 1 Introduction
Partial Differential Equations (PDEs) mathematically model a wide variety of interesting phenomena in physics and engineering, making their study a matter of great scientific importance [9]. Given the paucity of explicit solution formulas, numerical methods [34], such as finite difference, finite element and spectral methods constitute the main tools in the quantitative study of PDEs. Unfortunately, the computational cost of existing numerical methods or solvers for PDEs can be prohibitively expensive, particularly in the context of problems such as uncertainty quantification, inverse problems, control, and design, where PDE solvers have to be called multiple times [35]. This necessitates the design of fast and accurate surrogates for PDE solvers.
@@
In this context, machine learning (ML) and AI-inspired algorithms are increasingly being used as surrogates for PDE solvers, see [30] and the references therein. In particular, operator learning, i.e., where the task at hand is to learn the underlying solution operator of a PDE from data, has emerged as the dominant paradigm in the application of ML/AI techniques for PDEs. A large variety of operator learning algorithms have recently been proposed, with examples being DeepONets [7, 27], Fourier Neural Operator (FNO) [20], Low-rank Neural Operators [19], Convolutional Neural Operators (CNO) [37] and transformer-based operator learning algorithms [6, 33, 10, 12, 1], among many others.
An overwhelming majority of these operator learning algorithms or neural operators [15] assume that the underlying domain is Cartesian and is discretized on a regular grid; however, in practice, and particularly in engineering, PDEs are posed on domains with very complex and highly non-Cartesian geometries. Thus, the direct application of many of the aforementioned operator learning algorithms is difficult or impractical. Therefore, we require algorithms that can learn the solution operators of PDEs on arbitrary domains.
Various approaches have been proposed to address this question. To start, methods such as the FNO and the CNO, originally designed for Cartesian domains, can be extended to non-Cartesian domains via domain masking [37, 12]. Another approach for extending FNO to arbitrary domains is to directly evaluate the discrete Fourier transform with batched matrix multiplications rather than FFT [23]. Moreover, one can transform the underlying problem on a non-Cartesian domain to one on Cartesian domains either through extensions [14], learned diffeomorphisms as in Geo-FNO [21], or through learned interpolations between non-Cartesian physical domain and a Cartesian latent domain via graph neural operators [18] as in GINO [22] and references therein.
@@
A different approach to learning operators for PDEs on arbitrary domains is provided by graph neural networks or GNNs [5]. This is an ideal ML framework for handling relational data such as molecules, networks, or unstructured computational grids. The use of graph neural networks for simulating PDEs has received widespread attention [38, 39, 32, 4, 8], to name a few. However, genuine learning of operators, rather than a specific finite-dimensional discretization of them, requires the ability of the underlying algorithm to display some form of resolution invariance [15, 2], i.e., the operator learning algorithm should be able to process inputs and outputs on different spatial resolutions, both during training and testing, without any decrease in its overall accuracy. A priori, it is unclear if the existing GNNs for simulating PDEs on arbitrary domains satisfy this key property of resolution invariance.
Another key aspect of operator learning for PDEs is the issue of temporal discretization when approximating solution operators of time-dependent PDEs. These operators are continuous in time and the algorithm approximating them should be able to retain this continuity. However, existing frameworks for learning time-dependent PDEs are often based on auto-regressive rollouts [4, 38, 39, 8] and cannot be evaluated at arbitrary times without additional interpolations, limiting their applicability in the context of time-dependent PDEs.
Thus, there is a pressing need for an ML algorithm that can learn solution operators of PDEs on arbitrary domains, which can also deal with multiple spatial resolutions and provide appropriate temporal behavior when used for time-dependent PDEs, while simultaneously providing precise and accurate approximations to the underlying operator. The design of such an algorithm is the main point of this article.
To this end, we propose Region Interaction Graph Neural Operator (RIGNO) as an effective neural operator to learn PDEs on arbitrary domains. RIGNO uses GNNs end-to-end and does not require any transformations into Cartesian domains at any stage. It is constructed for multiscale information processing and many novel elements are introduced to enable RIGNO to provide spatial resolution invariance and to directly and continuously approximate the solution operator of time-dependent PDEs for any time. Through extensive numerical experiments, we demonstrate that RIGNO provides extremely accurate approximations to the solution operators of a variety of PDEs with domain geometries including Cartesian grids, structured and unstructred meshes, and random point clouds. Thus, we present a GNN-based framework for accurately and robustly learning the solution operators of PDEs on arbitrary domains.
@@ 2 Methods
Problem Formulation. We consider the following generic time-dependent PDE with $\Omega_{t}=(0,T)$ the time domain, $\Omega_{x}\subset\mathbb{R}^{d}$ the $d$-dimensional spatial domain, $u\in\mathcal{C}(\overline{\Omega}_{t};\mathcal{X})$ the solution of the PDE with $\mathcal{X}\subset L^{p}(\Omega_{x};\mathbb{R}^{s})$ for some $1\leq p<\infty$, $c\in\mathcal{Q}$ a spatially varying coefficient, $\mathcal{F}$ a differential operator, $\mathcal{B}$ a boundary operator, and $a\in\mathcal{X}$ the initial condition,
$\partial_{t}u=\mathcal{F}(c,t,u,\nabla_{x}u,\nabla^{2}_{x}u,\dots),\quad\forall(t,x)\in\Omega_{t}\times\Omega_{x},$ (1)
$\mathcal{B}u(t,x)=0,\quad\forall(t,x)\in\Omega_{t}\times\partial\Omega_{x},$
$u(0,x)=a(x),\quad\forall x\in\Omega_{x}.$
@@
As shown in Figure 1, our model is based on the well-known encode-process-decode paradigm [38, 39]. All the three components, i.e., encoder, processor, and decoder are graph neural networks (GNNs). The methods presented in this section are implemented using JAX [3] and are available on Github: https://github.com/camlab-ethz/rigno. As the aim is to learn the (extended) solution operator $\mathcal{G}^{\dagger}$ (2), the inputs to our model are, in general, i) a set of coordinates in the domain $\boldsymbol{\rm x}:=\{\boldsymbol{\rm x}_{i}\in\Omega_{x}\}_{i}$, where $\Omega_{x}$ is the spatial domain, ii) $\boldsymbol{u}:=\{\boldsymbol{u}_{i}\in\mathbb{R}^{s}\}_{i}$, discretized values of a function at known coordinates $\boldsymbol{\rm x}$, iii ) $\boldsymbol{c}:=\{\boldsymbol{c}_{i}\}_{i}$, discretized values of the spatial coefficient at $\boldsymbol{\rm x}$ coordinates, iv) the time $t$ and v) the lead time $\tau$. The model’s output $\boldsymbol{r}:=\{\boldsymbol{r}_{i}\in\mathbb{R}^{s}\}_{i}$ represent discretized values of a parameterized map at the coordinates $\boldsymbol{\rm x}$,
$\boldsymbol{r}=\mathcal{T}_{\theta}(\boldsymbol{u},\boldsymbol{\rm x},\boldsymbol{c},t,\tau).$ (3)
@@
The map $\mathcal{T}_{\theta}$ is a GNN and will process data that are defined on graphs. As shown in Figure 1 (for an idealized two-dimensional domain) and following ideas in [16], we consider two underlying sets of nodes. The first set of nodes corresponds to the unstructured point cloud on which the input (and output) data is defined. We call this point cloud as the physical mesh and the nodes that constitute it as physical nodes. We denote them as $\boldsymbol{V}_{\mathfrak{P}}=\{\boldsymbol{v}_{i}^{\mathfrak{P}}\}_{i}$. Note that we do not require any explicit relational information on these physical nodes.
Most of our processing will take place in another bespoken set of nodes that we call as regional nodes and the point cloud constituted by them as the regional mesh, denoted as $\boldsymbol{V}_{\mathfrak{R}}=\{\boldsymbol{v}_{i}^{\mathfrak{R}}\}_{i}$. Here, the regional nodes are constructed by randomly sub-sampling the set of physical nodes with a sampling ratio which is a hyperparameter of our model. A graph is constructed to connect the regional nodes (see Figure 1) and the resulting bidirectional regional mesh edges are denoted by $\boldsymbol{E}_{\mathfrak{R}\leftrightarrow\mathfrak{R}}=\{\boldsymbol{e}_{ij}^{\mathfrak{R}\leftrightarrow\mathfrak{R}}\}_{ij}$.
Moreover, we need to pass information between the physical nodes and the regional nodes, necessitating the construction of graphs between them. To this end, we connect physical nodes to regional nodes (Figure 1) by unidirectional edges $\boldsymbol{E}_{\mathfrak{P}\to\mathfrak{R}}=\{\boldsymbol{e}_{ij}^{\mathfrak{P}\to\mathfrak{R}}\}_{ij}$. These edges are constructed such that each regional node constitutes a sub-region for a subset of (nearby) physical nodes and serves to accumulate information from them. Similarly, information from the regional mesh can be transmitted to the physical mesh by constructing a graph with unidirectional edges $\boldsymbol{E}_{\mathfrak{R}\to\mathfrak{P}}=\{\boldsymbol{e}_{ij}^{\mathfrak{R}\to\mathfrak{P}}\}_{ij}$. As shown in Figure 1, these edges are constructed to accumulate information from multiple regional nodes into one physical node. Thus, construction of these unidirectional edges provides a very flexible upscaling and downscaling strategy for our model.
@@
Needless to say, when possible, the graph construction must be adapted to the boundary conditions of the PDE (1). We postpone the rather technical description of graph construction and incorporation of boundary conditions, particularly the novel additions to the model design to enforce periodic boundary conditions, to SM Section A.
Once the underlying graphs have been constructed, we must also provide geometric information to the GNNs. To this end, we impose translation equivariance by employing linear normalized coordinates $\boldsymbol{\zeta}_{i}\in[-1,+1]$ and angular normalized coordinates $\boldsymbol{\alpha}_{i}\in[0,2\pi]$ defined as
$\boldsymbol{\zeta}_{i}:=2\frac{\boldsymbol{\rm x}_{i}-\boldsymbol{\rm x}_{\rm c}}{\boldsymbol{L}}-1;\qquad\boldsymbol{\alpha}_{i}:=\pi(\boldsymbol{\zeta}_{i}+1).$
The structural features of an edge $\boldsymbol{e}_{ij}$ connecting node $i$ to node $j$ is $\boldsymbol{\varphi}_{ij}:=\left[\tilde{\boldsymbol{\zeta}}_{ij}\diamond\|\tilde{\boldsymbol{\zeta}}_{ij}\|_{2}\right]$; with $\tilde{\boldsymbol{\zeta}}_{ij}:=(\boldsymbol{\zeta}_{j}-\boldsymbol{\zeta}_{i})/(2\sqrt{d})$ being the relative coordinates of the connected nodes normalized by the diameter of the domain, $2\sqrt{d}$. We denote the vertical concatenation of column vectors by $\diamond$. The structural features of a node $\boldsymbol{v}_{i}$ with general boundary conditions are simply its normalized coordinates $\boldsymbol{\varphi}_{i}:=\boldsymbol{\zeta}_{i}$. With periodic boundary conditions (see SM Section B), we concatenate sines and cosines of angular normalized coordinates with $K_{\rm freq}$ frequencies:
$\boldsymbol{\varphi}_{i}=\left[\sin(k\boldsymbol{\alpha}_{i})\diamond\cos(k\boldsymbol{\alpha}_{i})\right],\quad 1\leq k\leq K_{\rm freq}.$ (4)
For regional nodes, the coverage radius of each node, $R_{i}$, is also concatenated with the other structural features. Once the inputs, graphs and geometric information are available, our model processes this information by employing the following constituent GNNs.
Encoder. Our encoder (Figure 1) is designed to transmit information from the underlying physical mesh (representing a discretization of the domain $\Omega_{x}$) to the (coarser) regional mesh, while at the same time, also embedding inputs into a latent space to enable representation learning. These are coupled into the following encoder module, consisting of two steps. In the first step, all the features are mapped into suitable latent spaces with,
[…]
Here and everywhere in this paper, FF denotes a feed-forward block (see Figure 1), which consists of a shallow multi-layer perceptron (MLP) followed by normalization layers. Next, the second step is a single message-passing block, with residual updates, to move the information from the physical mesh to the regional mesh:
Processor. The processor is designed to process the latent information on the regional mesh by $P$ sequential message-passing steps with residual updates and skip connections to allow for gradients to flow for large $P$, in order to facilitate stable training. The $p$-th message-passing step in the processor updates the regional nodes and the regional mesh edges as:
Decoder. The decoder serves to transfer processed information from the regional mesh back to the physical mesh (Figure 1). To this end, the first step in the decoder is a single message-passing block:
Finally, in the second step of the decoder, the physical node feature vectors are independently scaled down to the desired dimension to form the output:
Summarizing, as our model involves the learned interaction between nodes that each represent aggregated information from an underlying sub-region, we call it as Region Interaction Graph Neural Operator or RIGNO.
@@
Inspired by dropout [40], we employ random masking of edges in each message-passing step of RIGNO. Before each message-passing step (e.g., step $p$ of the processor (7)), each edge is masked with a given probability and does not participate in that specific message-passing block. In the next message-passing step (e.g., step $p+1$ of the processor), a new set of edges is randomly masked. Since the messages computed from the neighbors are averaged, there is no need for any further configuration to apply this at training or inference.
As shown in the next section, edge masking enables us to i) make RIGNO (approximately) resolution invariant, ii) reduce computational complexity by sparsely sampling edges during each step, and iii) provide a method for quantifying uncertainty in the predictions of RIGNO.
@@
We recall that our goal is to approximate the extended solution operator $\mathcal{G}^{\dagger}(u(t),c,t,\tau)$ (2), which is a function of the input time $t$ and the lead time $\tau$. Both of these scalars are provided to the encoder of RIGNO as inputs and embedded into suitable high-dimensional latent spaces. However, merely providing lead time as an input may not be sufficient to learn the solution operator in deep models. Hence, we introduce lead time as an additional conditioning in each hidden layer of RIGNO (see Figure 1). We do this by adapting the normalization layers employed within our model. To this end, let ${\bf y}$ be an output vector for any layer and normalize it by an appropriate normalization layer (BatchNorm, LayerNorm, InstanceNorm, etc.) to obtain the normalized output $\widetilde{{\bf y}}$. We further condition it with the lead time $\tau$ by setting,
$\hat{{\bf y}}=(1+\tau\gamma(\tau))\odot\widetilde{{\bf y}}+\tau\lambda(\tau),$ (10)
with $\gamma,\lambda$ being small learnable MLPs.
Finally, we need to provide a time stepping strategy to update our model in time. In the context, RIGNO can be employed in a very general time stepping mode by setting,
@@
where $\alpha,\beta$ are scalar parameters, $\mathcal{T}_{\theta}$ is the output of RIGNO, and $\mathcal{G}_{\theta}\approx\mathcal{G}^{\dagger}$ is the final approximation to the extended solution operator (2) of the PDE (1).
Setting $\alpha=0$ and $\beta=1$ directly recovers an approximation to the output of the extended solution operator. In this case, RIGNO can be interpreted to directly provide the solution of the PDE (1) at time $t+\tau$, given the solution at time $t$, coefficient $c$ and the lead time $\tau$ as inputs. Similarly, setting $\alpha=\beta=1$ yields the residual of the solution at the later time, given the solution at current time. On the other hand, setting $\alpha=1$ and $\beta=\tau$ leads to the model output approximating the time derivative of the solution operator as (11) is simply an approximation of the Taylor expansion in time of the solution operator, truncated at leading order. A clear advantage of this time-derivative update is that it automatically enforces time continuity by simply setting $\tau=0$. We will explore these choices for time updates in our numerical experiments to determine which ones are the most suitable.
@@
For simplicity of exposition here, we fix a final time $T$ and a fixed discretization $\Omega_{t}=\cup_{n=1}^{N}[t_{n-1},t_{n}]$. We assume that training samples are provided as trajectories $\{u_{m}(t_{n})\}_{n=0}^{N}$, for $1\leq m\leq M$, with each $u_{m}$ being the solution to the PDE (1) with initial condition $a_{m}$ and coefficient $c_{m}$. Given these training trajectories and the fact that our model can be evaluated at any time, there are multiple ways in which the RIGNO model, denoted here by $\mathcal{G}_{\theta}$ can be trained. We follow the recently proposed all2all training strategy [12] as training input-output pairs scale quadratically with respect to the number of time steps for time-dependent problems. Hence, the goal of training is to find parameters $\theta\in\Theta$ such that the following loss function can be be minimized,
${\mathcal{J}}:=\frac{1}{M\widehat{N}}\sum_{m,n,t_{k}\leq t_{n}}\mathcal{L}\left[u_{m}({t_{n}}),\;\widehat{u}_{m}^{t_{k}}({t_{n}})\right],$ (12)
$\widehat{u}_{m}^{t_{k}}({t_{n}})=\mathcal{G}_{\theta}\left(u_{m}({t_{k}}),c_{m},\;t_{k},\;t_{n}-t_{k})\right),$
@@
with $\widehat{N}=\frac{1}{2}(N+1)(N+2)$ and $\mathcal{L}$ being any measure of the mismatch between the ground truth and the model prediction, for instance either the mean absolute or the mean squared error.

### v2

@@ Abstract
Learning the solution operators of PDEs on arbitrary domains is challenging due to the diversity of possible domain shapes, in addition to the often intricate underlying physics. We propose an end-to-end graph neural network (GNN) based neural operator to learn PDE solution operators from data on point clouds in arbitrary domains. Our multi-scale model maps data between input/output point clouds by passing it through a downsampled regional mesh. The approach includes novel elements aimed at ensuring spatio-temporal resolution invariance. Our model, termed RIGNO, is tested on a challenging suite of benchmarks composed of various time-dependent and steady PDEs defined on a diverse set of domains. We demonstrate that RIGNO is significantly more accurate than neural operator baselines and robustly generalizes to unseen resolutions both in space and in time. Our code is publicly available at github.com/camlab-ethz/rigno.
@@ 1 Introduction
Partial Differential Equations (PDEs) mathematically model a wide variety of interesting phenomena in physics and engineering, making their study a matter of great scientific importance [9]. Given the paucity of explicit solution formulas, numerical methods [36], such as finite difference, finite element and spectral methods constitute the main tools in the quantitative study of PDEs. The computational cost of existing numerical methods or solvers for PDEs can be prohibitively expensive, particularly in the context of problems such as uncertainty quantification, inverse problems, control, and design, where PDE solvers have to be called multiple times [35]. This necessitates the design of fast and accurate surrogates for PDE solvers.
@@
In this context, machine learning (ML) algorithms are increasingly being used as surrogates for PDE solvers, see [31] and the references therein. In particular, operator learning, i.e., where the task at hand is to learn the underlying solution operator of a PDE from data, has emerged as the dominant paradigm in the application of ML techniques for PDEs. A large variety of operator learning algorithms have recently been proposed, with examples being DeepONets [7, 28], Fourier Neural Operator (FNO) [20], Low-rank Neural Operators [22], Convolutional Neural Operators (CNO) [38] and transformer-based operator learning algorithms [6, 34, 12, 13, 1], among many others. The success of these approaches may largely be contributed to their multi-scale learning. For example, the FNO learns frequencies of different scales within the data, CNO learns by down- and up-sampling via convolutions, and transformer-based algorithms often use multi-scale windowing. As a result, these algorithms learn both short- and long-range physical dependencies of the solutions of PDEs.
An overwhelming majority of these operator learning algorithms or neural operators [16] assume that the underlying domain is Cartesian and is discretized on a regular grid; however, in practice, and particularly in engineering, PDEs are posed on domains with very complex and highly non-Cartesian geometries. Thus, the direct application of many of the aforementioned operator learning algorithms is difficult or impractical. Therefore, we require algorithms that can learn the solution operators of PDEs on arbitrary domains. Various approaches have been proposed to extend neural operators to arbitrary domains. To start, methods such as the FNO and the CNO, originally designed for Cartesian domains, can be extended to non-Cartesian domains via domain masking [38, 13]. Another approach for extending FNO to arbitrary domains is to directly evaluate the discrete Fourier transform rather than FFT [24]. Moreover, one can transform the underlying problem to a Cartesian domain either through extensions [15], learned diffeomorphisms as in Geo-FNO [19], or through learned interpolations between non-Cartesian physical domain and a Cartesian latent domain via graph neural operators [21] as in GINO [23] and references therein. Applying attention mechanisms to a downsampled representation, as demonstrated in UPT [1] and Transolver [45], offers another viable approach. A conceptually distinct line of work leverages implicit neural representations [41, 42], which accommodate arbitrary geometries by encoding input and output functions into a shared latent vector that can be continuously queried at any spatial coordinate. The mapping between input and output functions is then learned in this latent space through a separate training stage.
@@
An alternative approach to learn operators for PDEs on arbitrary domains is provided by graph neural networks (GNNs) [5]. This is an ideal ML framework for handling relational data such as molecules, networks, or unstructured computational grids. The use of graph neural networks for simulating PDEs has received widespread attention [40, 39, 33, 4, 8], to name a few. Particularly, approaches such as MeshGraphNet [33] and its derivatives [10, 17] continue to demonstrate success in learning physical simulations. Graphs are able to implicitly capture the challenging collocation point distributions of unstructured meshes and arbitrary point clouds, making methods such as MeshGraphNet ideal for the non-Cartesian domain. However, genuine learning of operators, rather than a specific finite-dimensional discretization of them, requires the ability of the underlying algorithm to display some form of space-time resolution invariance [16, 2], i.e., the operator learning algorithm should be able to process inputs and outputs on different spatial and temporal resolutions, both during training and testing, without any decrease in its overall accuracy. Existing graph-based approaches such as MeshGraphNet cannot satisfy this key property, and even many neural operators on Cartesian grids do not inherently capture temporal resolution invariance. Moreover, due to the local nature of their graph construction, MeshGraphNets struggle to capture long-range dependencies within the data. Therefore, given the success of multi-scale neural operators such as FNO and CNO, the question naturally arises: does there exist a complementary, multi-scale, graph-based approach which is capable of learning space-time resolution invariant operators on complicated distributions of input points?
In this work, we answer this question affirmatively by proposing RIGNO, a Region Interacting Graph Neural Operator. RIGNO is an end-to-end graph neural network based on the encode-process-decode paradigm [40, 39], relying on an algorithm to construct a multi-scale regional mesh. We also incorporate several novel elements which, in combination with the algorithmic developments, push this algorithm to achieve state-of-the-art performance. The contributions of this work are as follows:
  • The introduction of a flexible regional mesh which simultaneously incorporates resolution invariance and multi-scale feature learning.
  • An edge masking technique for GNNs in the context of scientific machine learning, which improves resolution invariance, facilitates optimization during training, and can provide model uncertainty estimates at inference.
  • The introduction of temporal fractional pairing, a fine-tuning strategy for time-continuous neural operators that unlocks generalization to higher time resolutions.
  • Extensive experimental results that demonstrate significant improvements to the state-of-the-art across a variety of benchmarks with domain geometries including structured and unstructured meshes, random point clouds, and Cartesian grids.
@@ 2 Methods
Problem formulation.
We consider the following generic time-dependent PDE with $\Omega_{t}=(0,T)$ the time domain, $\Omega_{x}\subset\mathbb{R}^{d}$ the $d$-dimensional spatial domain, $u\in\mathcal{C}(\overline{\Omega}_{t};\mathcal{X})$ the solution of the PDE with $\mathcal{X}\subset L^{p}(\Omega_{x};\mathbb{R}^{s})$ for some $1\leq p<\infty$, $c\in\mathcal{Q}$ a spatially varying coefficient, $\mathcal{F}$ a differential operator, $\mathcal{B}$ a boundary operator, and $a\in\mathcal{X}$ the initial condition,
$\partial_{t}u=\mathcal{F}(c,t,u,\nabla_{x}u,\nabla^{2}_{x}u,\dots),$ $\forall(t,x)\in\Omega_{t}\times\Omega_{x},$ (1)
$\mathcal{B}u(t,x)=0,$ $\forall(t,x)\in\Omega_{t}\times\partial\Omega_{x},$
$u(0,x)=a(x),$ $\forall x\in\Omega_{x}.$
@@
As shown in Figure 1, our model is based on the well-known encode-process-decode paradigm [40, 39]. All the three components, i.e., encoder, processor, and decoder are graph neural networks (GNNs). The methods presented in this section are implemented using JAX [3] and are available on GitHub [^1] https://github.com/camlab-ethz/rigno. As the aim is to learn the (extended) solution operator $\mathcal{G}^{\dagger}$ (2), the inputs to our model are, in general, i) a set of coordinates in the domain $\boldsymbol{\rm x}:=\{\boldsymbol{\rm x}_{i}\in\Omega_{x}\}_{i}$, where $\Omega_{x}$ is the spatial domain, ii) $\boldsymbol{u}:=\{\boldsymbol{u}_{i}\in\mathbb{R}^{s}\}_{i}$, discretized values of a function at known coordinates $\boldsymbol{\rm x}$, iii ) $\boldsymbol{c}:=\{\boldsymbol{c}_{i}\}_{i}$, discretized values of the spatial coefficient at $\boldsymbol{\rm x}$ coordinates, iv) the time $t$ and v) the lead time $\tau$. The model’s output $\boldsymbol{r}:=\{\boldsymbol{r}_{i}\in\mathbb{R}^{s}\}_{i}$ represent discretized values of a parameterized map at the coordinates $\boldsymbol{\rm x}$ and is obtained by $\boldsymbol{r}=\mathcal{T}_{\theta}(\boldsymbol{u},\boldsymbol{\rm x},\boldsymbol{c},t,\tau)$.
@@
The map $\mathcal{T}_{\theta}$ is a GNN and will process data that are defined on graphs. As shown in Figure 1 (for an idealized two-dimensional domain) and following ideas in [17], we consider two underlying sets of nodes. The first set of nodes corresponds to the unstructured point cloud on which the input (and output) data is defined. We call this point cloud as the physical mesh and the nodes that constitute it as physical nodes. We do not perform any explicit information processing on these physical nodes. Most of our processing will take place in another bespoken set of nodes that we call as regional nodes and the point cloud constituted by them as the regional mesh. Here, the regional nodes are constructed by randomly sub-sampling the set of physical nodes with a sampling ratio which is a hyperparameter of our model. A graph is constructed to connect the regional nodes (see Figure 1) via bidirectional regional mesh edges. Moreover, we need to pass information between the physical nodes and the regional nodes, necessitating the construction of graphs between them. To this end, we connect physical nodes to regional nodes (Figure 1) by unidirectional edges. These edges are constructed such that each regional node constitutes a circular sub-region containing a subset of nearby physical nodes and serves to accumulate information from them. Based on the local density of regional nodes, a minimum support radius is calculated for each of them such that their union covers the whole domain. Information from the regional mesh can be transmitted to the physical mesh in a similar way by constructing a graph with unidirectional edges. As shown in Figure 1, these edges are constructed to accumulate information from multiple regional nodes into one physical node. Thus, construction of these unidirectional edges provides a very flexible upscaling and downscaling strategy for our model.
Needless to say, when possible, the graph construction must be adapted to the boundary conditions of the PDE (1). Once the underlying graphs have been constructed, we must also provide geometric information to the GNNs. We normalize the given coordinates and prepare them for all nodes and all edges accordingly. We postpone the rather technical description of graph construction, geometric features, and incorporation of boundary conditions, particularly the novel additions to the model design to enforce periodic boundary conditions, to SM Section A. Once the inputs, graphs and geometric information are available, the first step is to lift the initial features of all nodes and all edges to suitable high-dimensional latent spaces. To this end, feed-forward blocks are applied independently on the initial features of each node and edge. Each feed-forward block consists of a shallow multi-layer perceptron (MLP) followed by normalization layers (see Figure 1). Next, our model processes this high-dimensional latent information by employing the following constituent GNNs.
@@
Main components.
The encoder is designed to transmit information from the underlying physical mesh (representing a discretization of the domain $\Omega_{x}$) to the coarser regional mesh. This is done by performing a single message passing step with residual updates. The processor then operates on the latent information defined over the regional mesh by performing $P$ sequential message passing steps. We incorporate residual updates and skip connections to maintain gradient flow and promote stable training for large $P$. The decoder serves to transfer processed information from the regional mesh back to the physical mesh. The first step in the decoder is a single message-passing block. In the second step of the decoder, the physical node feature vectors are independently scaled down to the desired dimension to form the output $\boldsymbol{r}$. Summarizing, as our model involves the learned interaction between nodes that each represent aggregated information from an underlying sub-region, we call it as Region Interaction Graph Neural Operator or RIGNO. More details about the GNN components are provided in SM B.
@@
Considering the graph structure in the encoder as an example, each regional node aggregates information from multiple physical nodes. For high-resolution inputs, many of these physical nodes can be redundant, as the underlying function can often be accurately reconstructed (e.g., through interpolation) using a smaller subset of nodes. During training, the model may implicitly disregard some of these redundant nodes and instead focus on a subset of dominant or central connections. However, at test time, when the input resolution differs, these dominant nodes may no longer exist, resulting in degraded performance. We identify this phenomenon as a primary cause of the resolution sensitivity commonly observed in GNN-based operator learning models. To address this, we use random masking of edges in all message-passing blocks of RIGNO. Before each message passing step (e.g., step $p$ of the processor), each edge is masked with a given probability and does not participate in that specific message-passing block. In the next message passing step (e.g., step $p+1$ of the processor), a different set of edges is randomly masked. Through random masking of the edges during training, the model cannot consistently rely on the same high-correlation connections. In the absence of these edges, the model is forced to use the weaker correlations. Additionally, learning from weaker or secondary correlations in the processor enables the model to better integrate information across multiple spatial scales, rather than overfitting to dominant patterns in the training resolution. Although loosely inspired by dropout [43], edge masking is conceptually very different from it. While dropout masks trainable parameters of MLPs, edge masking disables pairwise interaction of spatial entities (nodes). This allows the model to discover weaker spatial correlations in the data. The set of the masked edges randomly changes in every epoch, which allows a single message-passing block to learn from all the edges during training. Since the messages computed from the neighbors are averaged, there is no need for any further configuration at training or inference. It is also possible to use a different masking probability or to completely disable it at any point during training or inference. As shown in the next section, edge masking enables us to i) make RIGNO (approximately) resolution invariant, ii) obtain significantly lower errors, iii) reduce computational complexity by sparsely sampling edges during each step, and iv) provide a method for quantifying uncertainty in the predictions of RIGNO.
@@
We recall that our goal is to approximate the extended solution operator $\mathcal{G}^{\dagger}(u(t),c,t,\tau)$ (2), which is a function of the input time $t$ and the lead time $\tau$. Both of these scalars are provided to the encoder of RIGNO as inputs and embedded into suitable high-dimensional latent spaces. However, merely providing lead time as an input may not be sufficient to learn the solution operator in deep models. Following [13], we introduce lead time as an additional conditioning in each hidden layer of RIGNO (see Figure 1). We do this by adapting the normalization layers employed within our model. Let $\bf y$} be an output vector for any layer and normalize it by an appropriate normalization layer (e.g., BatchNorm, LayerNorm) to obtain the normalized output $\widetilde{{\bf y}}$. We further condition $\widetilde{{\bf y}}$ with the lead time $\tau$ by setting $\hat{{\bf y}}=(1+\tau\gamma(\tau))\odot\widetilde{{\bf y}}+\tau\lambda(\tau)$, with $\gamma,\lambda$ being small learnable MLPs.
Finally, we need to provide a time marching strategy to update our model in time. In this context, RIGNO can be employed in a very general time marching mode by setting,
@@
where $\alpha,\beta$ are scalar parameters, $\mathcal{T}_{\theta}$ is the output of RIGNO, and $\mathcal{G}_{\theta}\approx\mathcal{G}^{\dagger}$ is the final approximation to the extended solution operator (2) of the PDE (1). Setting $\alpha=0$ and $\beta=1$ directly recovers an approximation to the output of the extended solution operator. In this case, RIGNO can be interpreted to directly provide the solution of the PDE (1) at time $t+\tau$, given the solution at time $t$, coefficient $c$ and the lead time $\tau$ as inputs. Similarly, setting $\alpha=\beta=1$ yields the residual of the solution at the later time, given the solution at current time. On the other hand, setting $\alpha=1$ and $\beta=\tau$ leads to the model output $\mathcal{T}_{\theta}$ approximating the time derivative of the solution operator as (3) is simply an approximation of the Taylor expansion in time of the solution operator, truncated at leading order. A clear advantage of this time-derivative update is that it automatically enforces consistency for $\tau=0$. We will explore these choices for time updates in our numerical experiments to determine which ones are the most suitable.
@@
For simplicity of exposition, here, we fix a final time $T$ and consider a fixed discretization $\Omega_{t}=\cup_{n=1}^{N}[t_{n-1},t_{n}]$. We assume that training samples are provided as trajectories $\{u_{m}(t_{n})\}_{n=0}^{N}$, for $1\leq m\leq M$, with each $u_{m}$ being the solution to the PDE (1) with initial condition $a_{m}$ and coefficient $c_{m}$. Given these training trajectories and the fact that our model can be evaluated at any time, there are multiple ways in which the RIGNO model, denoted here by $\mathcal{G}_{\theta}$ can be trained. We follow the recently proposed all2all training strategy [13] as training input-output pairs scale quadratically with respect to the number of time steps for time-dependent problems. Hence, the goal of training is to find parameters $\theta\in\Theta$ such that the following loss function can be be minimized,
${\mathcal{J}}=\frac{1}{M\widehat{N}}\sum_{m,n,t_{k}\leq t_{n}}\mathcal{L}\left[u_{m}({t_{n}}),\;\widehat{u}_{m}^{t_{k}}({t_{n}})\right],\qquad\widehat{u}_{m}^{t_{k}}({t_{n}})=\mathcal{G}_{\theta}\left(u_{m}({t_{k}}),c_{m},\;t_{k},\;t_{n}-t_{k})\right),$ (4)
@@
with $\widehat{N}=\frac{1}{2}(N+1)(N+2)$ and $\mathcal{L}$ being a measure of the mismatch between the ground truth and the model prediction, for instance either the mean absolute or the mean squared error.